Exercise 1. The director of a state agency believes that the average starting salary for clerical employees in the state is less than $30,000 per year. To test her hypothesis, she has collected a simple random sample of 100 starting clerical salaries from across the state and found that the sample mean is $29,750.

  1. State the appropriate null and alternative hypotheses.

  2. Assuming the population standard deviation is known to be $2,500 and the significance level for the test is to be 0.05, what is the critical value (stated in dollars)?

  3. Referring to your answer in part b, what conclusion should be reached with respect to the null hypothesis?

  4. Referring to your answer in part c, which of the two statistical errors might have been made in this case? Explain.
Answer #1

(a)

Question:

State the appropriate null and alternative hypotheses.

Answer:
H0:Null Hypothesis:\mu \geq 30,000 ( the average starting salary for clerical employees in the state is not less than $30,000 per year.)

HA:Alternative Hypothesis:\mu < 30,000 ( the average starting salary for clerical employees in the state is less than $30,000 per year.) (Claim)

(b)

Question:

Assuming the population standard deviation is known to be $2,500 and the significance level for the test is to be 0.05, what is the critical value (stated in dollars)?

Answer:

\alpha =0.05

One Tail - Left Side Test

From Table, critical value of ZC = - 1.64

(c)

Question:

Referring to your answer in part b, what conclusion should be reached with respect to the null hypothesis?

Answer:
Test Statistic is given by:

Z=\frac{\bar{x}-\mu _{0}}{\sigma /\sqrt{n}}=\frac{29750-30000}{2500/\sqrt{100}}=-1.00

Since calculated value of Z = - 1.00 is greater than critical value of Z = - 1.64, the difference is not significant. Fail to reject null hypothesis.

Conclusion:
The data do not support the claim that the average starting salary for clerical employees in the state is less than $30,000 per year.

(d)

Question:

Referring to your answer in part c, which of the two statistical errors might have been made in this case? Explain.

Answer:
Since we have failed to reject null hypothesis, we might have made Type II Error because Type II Error is Failure to reject a false null hypothesis.